OpenStudy (anonymous):

solution 2x+3y=4 and -2-3y=-4

OpenStudy (anonymous):

add the equation to eliminate y: 2x - 2 = 0 x = 1 plug this into first equation: 2 + 3y = 4 3y = 2 y = 2/3 solution set is{1 , 2/3}

OpenStudy (mertsj):